# WavCraft: Audio Editing and Generation with Natural Language Prompts

### v2

@@ Abstract
We introduce WavCraft, a collective system that leverages large language models (LLMs) to connect diverse task-specific models for audio content creation and editing. Specifically, WavCraft describes the content of raw sound materials in natural language and prompts the LLM conditioned on audio descriptions and users’ requests. WavCraft leverages the in-context learning ability of the LLM to decomposes users’ instructions into several tasks and tackle each task collaboratively with audio expert modules. Through task decomposition along with a set of task-specific models, WavCraft follows the input instruction to create or edit audio content with more details and rationales, facilitating users’ control. In addition, WavCraft is able to cooperate with users via dialogue interaction and even produce the audio content without explicit user commands. Experiments demonstrate that WavCraft yields a better performance than existing methods, especially when adjusting the local regions of audio clips. Moreover, WavCraft can follow complex instructions to edit and even create audio content on the top of input recordings, facilitating audio producers in a broader range of applications. Our implementation and demos are available at https://github.com/JinhuaLiang/WavCraft.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT (OpenAI, 2023) have remarkably promoted the development of artificial intelligence-generated content (AIGC). Driven by large-scale pre-training on massive high-quality textual tokens and reinforcement learning from human feedback (RLHF), LLMs demonstrate advanced capacity in language analysis, rationale, and interaction. While LLMs have attracted increasing amount of attention on topics such as chain-of-thought (Wei et al., 2023a), interpretability (Zhao et al., 2024), and in-context learning (Wei et al., 2023b), they are limited to textual data and fail to engage with a broader range of AIGC tasks.
@@
AI-empowered agents have been devised to tackle more practical applications by equipping LLMs with task-specific modules (Qian et al., 2023). These agents (Shen et al., 2023; Huang et al., 2023) use the LLM to interpret a user query to some basic tasks and call task-specific modules (namely expert models) with an appropriate order. By using a modular approach, AI-driven agents are capable of solving intricate tasks without the requirement of additional training. In the audio domain, WavJourney (Liu et al., 2023e) proposed an AI-driven agent that synthesises an audio clip by connecting speech, audio, and music generative models. An audio script is created based on user instructions and complied into an executable computer program. The computer program then instructs various audio generative models to synthesize a recording. Despite its success, the current audio agents cannot use audio clips as input, hindering themselves from a broader range of audio generation applications. Considering the collaborative ability of LLMs and real-world need for multimodal interactive creation, a natural question arises: can we improve audio agents with the ability of audio analysis and transformation?
@@
In this work, we introduce WavCraft, an AI-empowered audio agent that leverages LLMs, together with a variety of expert models, to create audio content based on human instructions and available audio materials. Specifically, WavCraft describes the content of input audio clips with an audio analysis module. The audio description and the user query are wrapped up with a pre-defined instruction template and directed to an audio programmer module. The audio programmer applies an LLM to break down a complex audio content creation task into several basic ones, and generate an executable program to invoke modules like audio expert models, DSP functions, or logic operations. Following this modular approach, WavCraft is able to assemble a variety of audio creation tools with great flexibility. The overall framework of WavCraft and its featured use cases are shown in Figure 1.
@@
Features of WavCraft includes: 1) Adjustable. WavCraft can take available audio clips as raw materials and create audio content based on both user instruction and input audio. Compared with existing audio agents (Liu et al., 2023e), WavCraft is capable of a broader range of audio content creation, such as audio editing; 2) Modular. WavCraft can break down a comprehensive instruction into several basic audio tasks and thus handle a wide range of audio content generation tasks. In addition, the decompositional framework of WavCraft presents an explicit pathway of content creation, enhancing the explainability in the eyes of users; 3) Interactive. Exploiting the language analysis ability of the LLM, WavCraft interacts with users in multiple dialogues. During multi-round co-creation, the generated audio clips stay consistent with each other; 4) Creative. Based on the analysis of audio content and user’s blueprint, WavCraft leverages the LLM to narrate a story, infers instructions for expert models, and creates the audio content that fulfill the storyline. We refer to such ability of WavCraft to generate audio content without explicit user instruction as audio scriptwriting.
@@ 2 Related work
Language models for multi-modal tasks. Language models such as ChatGPT (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have achieved considerable progress in natural language processing. These models, featuring billions of parameters, are trained with massive high-quality training data to handle a variety of text-related tasks with a single model. To extend the open-world knowledge to more domains, subsequent works (Alayrac et al., 2022; Li et al., 2023) aligned audio, video, and/or image to text and proposed multi-modal foundational models in their respective domains. While multi-modal language models achieved state-of-the-art performance in downstream tasks, most of them are restricted to common text-guided generation tasks, such as text-to-audio synthesis (Liu et al., 2023b), where the model is not required to analyse complex user instructions.
@@
LLM-based agents. More recently, LLM-based agents have attracted an increasing amount of attention to tackle challenging, intricate applications by integrating language models with a set of task-specific models. Toolformer (Schick et al., 2023) was trained to decide when and where an API should be called and how to assemble outputs from different APIs. HuggingGPT (Shen et al., 2023) applied ChatGPT (OpenAI, 2023) as a controller to allocate existing neural networks in Huggingface (Shen et al., 2023). HuggingGPT is thus capable of solving diverse AI tasks across natural language, visual, and audio domains. Meanwhile, ViperGPT (Surís et al., 2023), VisProg (Gupta & Kembhavi, 2023), and RVP (Ge et al., 2023) have demonstrated the promise of visual agents on image/video analysis tasks. LLaVA-Plus (Liu et al., 2023c) extended the input query to visual domain by replacing LLM with visual language model (VLM). In the audio domain, AudioGPT (Huang et al., 2023) connected multiple audio neural networks and used ChatGPT to classify the user query into a predefined task. WavJourney (Liu et al., 2023e) used LLM to screenwrite the audio scripts and then generate audio clips by calling diverse audio generative models. Although considerable progress has been made in previous works to extend open-world knowledge in language models to multiple modalities, few of them can be prompted by non-text inputs, restricting their ability to many practical applications, especially audio editing.
@@
Audio Creation and Editing. Audio creation and editing are challenging parts of generative AI since they require models to not only understand the audio content but also modify the audio conditioned on input instructions. With the development of deep learning, generative models have demonstrated remarkable capacities to synthesise speech (Wang et al., 2023a), audio (Liu et al., 2023b; Kreuk et al., 2023; Borsos et al., 2022; Team et al., ), and music (Copet et al., 2023; Agostinelli et al., 2023). Existing audio generation methods are mainly dedicated to synthesising audio conditioned on different types of prompts, such as text description, voice style, and music melody. However, these methods are trained to generate audio from scratch and thus are ill-suited to editing tasks on existing audio. Recently, AUDIT (Wang et al., 2023b) was proposed to learn an end-to-end diffusion model to modify the audio content based on both text instructions and input audio. While AUDIT is capable of various basic editing tasks, including adding, removal, replacement, super-resolution, and infilling, it suffers from two drawbacks: 1) it cannot address complex editing tasks that combines these basic tasks; and 2) it cannot perform local changes on designated audio regions, limiting its application in real-world scenario.
@@
Task Input Output API Name Model Name
@@
Text-to-Speech Text Audio TTS Bark
Text-guided Source Separation Text, Audio Audio TSS AudioSep (Liu et al., 2023d)
Extract Audio Audio EXTRACT AudioSep (Liu et al., 2023d)
@@
Super resolution Audio Audio SR AudioSR
Drop Text Audio DROP AudioSep (Liu et al., 2023d)
Inpaint Audio Audio INPANT AudioLDM (Liu et al., 2023b)
@@ 3.1 Overall framework
Task decomposition. The audio programmer module in WavCraft then drives the LLM to generate an executable script conditioned on both the user query and the content of input audio clips. Specifically, WavCraft fills a pre-defined template with the input query and audio description and then directs the instructions to the LLM Compared with other AI programmers (Gupta & Kembhavi, 2023), WavCraft generates not only the code but also the comment for each line and the audio script. We found that these comments and the audio script facilitate the audio programmer module to generate code step by step, leading to high-quality output and explainable operations.
Code execution. WavCraft executes the generated scripts by calling a set of audio expert models. Table 1 lists the APIs constituting WavCraft. WavCraft consists of various publicly-available expert models: AudioGen (Kreuk et al., 2023) was used for text-to-audio generation; MusicGen (Copet et al., 2023) was adopted to music synthesis due to its high-fidelity performance based on text and/or melody. For text-to-speech generation, we use Bark [^1] https://huggingface.co/spaces/suno/bark, a state-of-the-art model that generates matched speech conditioned on the tone, pitch, emotion, and prosody of a given voice preset. For text-guided source separation, AudioSep (Liu et al., 2023d) is used to separate targeted sound tracks conditioned on language queries. AudioSR (Liu et al., 2023a) and AudioLDM (Liu et al., 2023b) are used for super-resolution and audio infilling, respectively. In addition, a series of DSP modules are introduced as well. We implement the DSP modules by using torchaudio (Yang et al., 2022) and audiomentations  [^2] https://github.com/iver56/audiomentations. It is noteworthy that these task-specific modules can be easily replaced with the alternative architectures.
@@ 4 Tasks
WavCraft provides a flexible framework that can address a diverse range of audio generative and editing tasks. We evaluate WavCraft on 5 basic tasks, involving adding, removal, replacement, super-resolution, and infilling. We also assess the advanced features of WavCraft on complex tasks.
@@
Removal. Given a mixture $M$ and one of its sound track $A$, the model is required to output a new audio clip $B$ by removing $A$ from $M$. Suppose $C_{A}$ and $C_{M}$ are the caption of A, respectively, an example of the instruction can be “Remove $C_{A}$ from $C_{M}$”.
Replacement. Given an audio clip $A$, a mixture $M$ and one of its sound track $B$, the model is required to output a new audio clip $C$ by replacing $B$ with $A$ in the same time slot. An example of the instruction can be “Replace $C_{B}$ with $C_{A}$”.
@@
Audio infilling. Given an audio clip $A$ where some parts are randomly masked, the model is required to complete the audio by filling the masked areas. An example of the instruction can be “Inpaint $A$”.
@@
Task AUDIT (Wang et al., 2023b) WavCraft
@@ 5.1 Experiments setup
To build up WavCraft, we used the GPT-4 model (OpenAI, 2023) as audio programming module and LTU (Gong et al., 2023) for audio analysis. We applied the publicly available models as audio expert models (shown in Table 1). We use 16 kHz sampling rate throughout the pipeline of audio editing and generation in line with the sampling rate of many integrated generative models (Kreuk et al., 2023; Copet et al., 2023). For the volume control of the generated audio content, we adopt the Loudness Unit Full Scale (LUFS) standard (International Telecommunication Union, 2020).
We evaluated WavCraft on audio editing and generation tasks separately. We compared WavCraft with AUDIT (Wang et al., 2023b), an state-of-the-art audio editing model, on diverse downstream tasks. For text-to-audio generation, we used AudioLDM (Liu et al., 2023b) and WavJourney (Liu et al., 2023e) for comparison. It is noteworthy that while WavJourney is also an LLM-based agent for audio content generation, it cannot take waveforms as inputs.
@@ 5.2 Evaluation metrics
For objective evaluation, we follows the evaluation protocols of existing audio generative models (Liu et al., 2023b; e; Wang et al., 2023b) to calculate several measurements: Frechet Audio Distance (FAD), Kullback-Leibler Divergence (KL), Inception Score (IS) and Log Spectral Distance (LSD) for evaluation. FAD measures the Frechet distance between reference and generated audio distributions of the embeddings extracted by a pre-trained VGGish model (Gemmeke et al., 2017). KL computes the similarity between logit distributions of two audio groups by using an audio tagging model, namely Patch-out Transformer (Koutini et al., 2022). IS reflects the variety and diversity of the generated audio group. Log Spectral Distance (LSD) calculates the distance between frequency spectrograms of output samples and target samples. A lower score of FAD, KL, or LSD indicates a better audio fidelity while a higher IS indicates a more diverse audio group (and thus more desirable for generated audio).

### v3

# WavCraft: Audio Editing and Generation with Large Language Models

@@ Abstract
We introduce WavCraft, a collective system that leverages large language models (LLMs) to connect diverse task-specific models for audio content creation and editing. Specifically, WavCraft describes the content of raw audio materials in natural language and prompts the LLM conditioned on audio descriptions and user requests. WavCraft leverages the in-context learning ability of the LLM to decomposes users’ instructions into several tasks and tackle each task collaboratively with the particular module. Through task decomposition along with a set of task-specific models, WavCraft follows the input instruction to create or edit audio content with more details and rationales, facilitating user control. In addition, WavCraft is able to cooperate with users via dialogue interaction and even produce the audio content without explicit user commands. Experiments demonstrate that WavCraft yields a better performance than existing methods, especially when adjusting the local regions of audio clips. Moreover, WavCraft can follow complex instructions to edit and create audio content on the top of input recordings, facilitating audio producers in a broader range of applications. Our implementation and demos are available at this https://github.com/JinhuaLiang/WavCraft.
@@ 1 Introduction
Large language models (LLMs) such as ChatGPT (OpenAI, 2023) have remarkably promoted the development of artificial intelligence-generated content (AIGC). Driven by large-scale pre-training on massive high-quality textual tokens and reinforcement learning from human feedback (RLHF), LLMs demonstrate advanced capacity in language analysis, rationale, and interaction. While LLMs have attracted increasing amount of attention on topics such as chain-of-thought (Wei et al., 2022), interpretability (Zhao et al., 2024), and in-context learning (Wei et al., 2023), they are limited to textual data and fail to engage with a broader range of AIGC tasks.
@@
AI-empowered agents have been devised to tackle more practical applications by equipping LLMs with task-specific modules (Qian et al., 2023). These agents (Shen et al., 2023; Huang et al., 2024) use the LLM to interpret a user query to some basic tasks and call task-specific modules (namely expert models) with an appropriate order. By using a modular approach, AI-driven agents are capable of solving intricate tasks without the requirement of additional training. In the audio domain, WavJourney (Liu et al., 2023d) proposed an AI-driven agent that synthesises an audio clip by connecting speech, audio, and music generative models. An audio script is created based on user instructions and compiled into an executable computer program. The computer program then invokes various audio generative models to synthesize a recording. Despite its success, the current audio agents cannot use audio clips as input, hindering themselves from a broader range of audio generation applications. Considering the collaborative ability of LLMs and the real-world need for multimodal interactive creation, a natural question arises: can we improve audio agents with the ability of audio analysis and transformation?
@@
In this work, we introduce WavCraft, an AI-empowered audio agent that leverages LLMs, together with a variety of expert models, to edit and generate audio content based on human instructions and available audio materials. Given a few audio references, WavCraft analyses the content of audio references with an audio analysis module to produce audio description. The audio description and the user query are wrapped up with a pre-defined instruction template and directed to an audio programmer module. The audio programmer applies an LLM to break down a complex audio content creation task into several basic ones, and generates an executable program to invoke modules like audio expert models, DSP functions, or logic operations. Following this modular approach, WavCraft is able to assemble a variety of audio creation tools with great flexibility. The overall framework of WavCraft and its featured use cases are shown in Figure 1.
@@
Features of WavCraft include: 1) Adjustable: WavCraft can take available audio clips as raw materials and create audio content based on both user instruction and input audio. Compared with existing audio agents (Liu et al., 2023d), WavCraft is capable of a broader range of audio content creation, such as sound infilling and removal; 2) Modular: WavCraft can break down a comprehensive instruction into several basic audio tasks and thus handle a wide range of audio content generation tasks. In addition, the decompositional framework of WavCraft presents an explicit pathway of content creation, enhancing the explainability in the eyes of users; 3) Interactive: Exploiting the language analysis ability of the LLM, WavCraft interacts with users in multiple dialogues. During multi-round co-creation, the generated audio clips stay consistent with each other; 4) Creative: Based on the analysis of audio content and user’s blueprint, WavCraft leverages the LLM to narrate a story, infers instructions for expert models, and creates the audio content that fulfill the storyline. We refer to such ability of WavCraft to generate audio content without explicit user instruction as audio scriptwriting.
@@ 2 Related work
Language models for multi-modal tasks. Language models such as ChatGPT (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have achieved considerable progress in natural language processing. These models, featuring billions of parameters, are trained with massive high-quality training data to handle a variety of text-related tasks with a single model. To extend the open-world knowledge to more domains, subsequent works (Alayrac et al., 2022; Li et al., 2023) have aligned audio, video, and/or image to text and proposed multi-modal foundational models in their respective domains. While multi-modal language models achieved state-of-the-art performance in downstream tasks, most of them are restricted to common text-guided generation tasks, such as text-to-audio synthesis (Liu et al., 2023a), where the model is not required to analyse complex user instructions.
@@
LLM-based agents. More recently, LLM-based agents have attracted an increasing amount of attention to tackle challenging, intricate applications by integrating language models with a set of task-specific models. Toolformer (Schick et al., 2023) was trained to decide when and where an API should be called and how to assemble outputs from different APIs. HuggingGPT (Shen et al., 2023) applied ChatGPT (OpenAI, 2023) as a controller to allocate existing neural networks in Huggingface (Shen et al., 2023). HuggingGPT is thus capable of solving diverse AI tasks across natural language, visual, and audio domains. Meanwhile, ViperGPT (Surís et al., 2023), VisProg (Gupta & Kembhavi, 2023), and RVP (Ge et al., 2023) have demonstrated the promise of visual agents on image/video analysis tasks. LLaVA-Plus (Liu et al., 2023b) extended the input query to visual domain by replacing LLM with visual language model (VLM). In the audio domain, AudioGPT (Huang et al., 2024) connected multiple audio neural networks and used ChatGPT to classify the user query into a predefined task. WavJourney (Liu et al., 2023d) used an LLM to screenwrite the audio scripts and then generate audio clips by calling diverse audio generative models. Although considerable progress has been made in previous works to extend open-world knowledge in language models to multiple modalities, few of them can be prompted by non-text inputs, restricting their ability to many practical applications, especially audio editing.
@@
Audio Creation and Editing. Audio creation and editing are challenging parts of generative AI since they require models to not only understand the audio content but also modify the audio conditioned on input instructions. With the development of deep learning, generative models have demonstrated remarkable capacities to synthesise speech (Wang et al., 2023), audio (Liu et al., 2023a; Kreuk et al., 2023; Borsos et al., 2023; Vyas et al., 2023), and music (Copet et al., 2023; Agostinelli et al., 2023). Existing audio generation methods are mainly dedicated to synthesising audio conditioned on different types of prompts, such as text description, voice style, and music melody. However, these methods are trained to generate audio from scratch and thus are ill-suited to editing tasks on existing audio. Recently, AUDIT (Wang et al., ) was proposed to learn an end-to-end diffusion model to modify the audio content based on both text instructions and input audio. While AUDIT is capable of various basic editing tasks, including adding, removal, replacement, super-resolution, and infilling, it suffers from two drawbacks: 1) it does not perform well in complex editing tasks that combines these basic tasks; and 2) it cannot perform local changes on designated audio regions, limiting its application in real-world scenario.
@@
{tblr}
width=cell21=c=5h, cell111=c=5h, hline1-2,11,21=-, Task & Input Output API Name Model name
@@
Text-to-Speech  Text Audio TTS Bark  [^1] https://github.com/suno-ai/bark
Text-guided Source Separation Text, Audio Audio TSS AudioSep (Liu et al., 2023c)
Extract Audio Audio EXTRACT AudioSep (Liu et al., 2023c)
@@
Super resolution Audio Audio SR AudioSR (Liu et al., 2024)
Drop Text Audio DROP AudioSep (Liu et al., 2023c)
Inpaint Audio Audio INPANT AudioLDM (Liu et al., 2023a)
@@ 3.1 Overall framework
Task decomposition. The audio programmer module in WavCraft then drives the LLM to generate an executable script conditioned on both the user query and the content of input audio clips. Specifically, WavCraft fills a pre-defined template with the input query and audio description and then directs the instructions to the LLM (see more in the Appendix A). Compared with other AI programmers (Gupta & Kembhavi, 2023), WavCraft generates not only the code but also the comment for each line and the audio script. We found that these comments and the audio script facilitate the audio programmer module to generate code step by step, leading to high-quality output and explainable operations.
Code execution. WavCraft executes the generated scripts by calling a set of audio expert models. Table 2 lists the APIs constituting WavCraft. WavCraft consists of various publicly-available expert models: AudioGen (Kreuk et al., 2023) was used for text-to-audio generation; MusicGen (Copet et al., 2023) was adopted to music synthesis due to its high-fidelity performance based on text and/or melody. For text-to-speech generation, we use Bark [^2] https://huggingface.co/spaces/suno/bark, a state-of-the-art model that generates matched speech conditioned on the tone, pitch, emotion, and prosody of a given voice preset. For text-guided source separation, AudioSep (Liu et al., 2023c) is used to separate targeted sound tracks conditioned on language queries. AudioSR (Liu et al., 2024) and AudioLDM (Liu et al., 2023a) are used for super-resolution and audio infilling, respectively. In addition, a series of DSP modules are introduced as well. We implement the DSP modules by using torchaudio (Yang et al., 2022) and audiomentations  [^3] https://github.com/iver56/audiomentations. It is noteworthy that these task-specific modules can be easily replaced with the alternative architectures.
@@ 4 Tasks
WavCraft provides a flexible framework that can address a diverse range of audio generative and editing tasks. We evaluate WavCraft on five basic tasks, involving adding, removal, replacement, super-resolution, and infilling. We also assess the advanced features of WavCraft on complex tasks, such as controllable editing and audio scriptwriting.
@@
Removal. Given a mixture $M$ and one of its sound tracks $A$, the model is required to output a new audio clip $B$ by removing $A$ from $M$. Suppose $C_{A}$ and $C_{M}$ are the captions of A, respectively, an example of the instruction can be “Remove $C_{A}$ from $C_{M}$”.
Replacement. Given an audio clip $A$, a mixture $M$ and one of its sound tracks $B$, the model is required to output a new audio clip $C$ by replacing $B$ with $A$ in the same time slot. An example of the instruction can be “Replace $C_{B}$ with $C_{A}$”.
@@
Audio infilling. Given an audio clip $A$ where some parts are masked, the model is required to complete the audio by filling the masked areas. An example of the instruction can be “Inpaint $A$”.
@@
Task AUDIT (Wang et al., ) WavCraft
@@ 5.1 Experiments setup
To build up WavCraft, we used the GPT-4 model (OpenAI, 2023) as audio programming module and LTU (Gong et al., 2024) for audio analysis. We applied the publicly available models as audio expert models (shown in Table 2). We use 16 kHz sampling rate throughout the pipeline of audio editing and generation in line with the sampling rate of many integrated generative models (Kreuk et al., 2023; Copet et al., 2023). For the volume control of the generated audio content, we adopt the Loudness Unit Full Scale (LUFS) standard (International Telecommunication Union, 2020).
We evaluated WavCraft on audio editing and generation tasks separately. We compared WavCraft with AUDIT (Wang et al., ), an state-of-the-art audio editing model, on diverse downstream tasks. For text-to-audio generation, we used AudioLDM (Liu et al., 2023a) and WavJourney (Liu et al., 2023d) for comparison. While WavJourney is also an LLM-based agent for audio content generation, it cannot take waveforms as inputs. We evaluate editing and text-to-audio generation task on AudioCaps (Kim et al., 2019) datasets. For editing tasks, we synthesis the database in according to 4 while directly using the evaluation split for the text-to-audio generation task. Please note that we will refer to the synthesised audio samples as the ground truth in the following sessions.
@@ 5.2 Evaluation metrics
For objective evaluation, we follows the evaluation protocols of existing audio generative models (Liu et al., 2023a; d; Wang et al., ) to calculate several measurements: Frechet Audio Distance (FAD), Kullback-Leibler Divergence (KL), Inception Score (IS) and Log Spectral Distance (LSD) for evaluation. FAD measures the Frechet distance between reference and generated audio distributions of the embeddings extracted by a pre-trained VGGish model (Gemmeke et al., 2017). KL computes the similarity between logit distributions of two audio groups by using an audio tagging model, namely Patch-out Transformer (Koutini et al., 2022). IS reflects the variety and diversity of the generated audio group. Log Spectral Distance (LSD) calculates the distance between frequency spectrograms of output samples and target samples. A lower score of FAD, KL, or LSD indicates a better audio fidelity while a higher IS indicates a more diverse audio group (and thus more desirable for generated audio). Subjective evaluation were carried out by Amazon Mechanical Turk  [^4] https://requester.mturk.com. We offered raters with detailed instructions and illustrative examples to ensure a thorough evaluation process. Each audio sample was rated from one to five by a minimum of 15 different raters. We collected the feedback from rates and calculate mean opinion score (MOS) to reflect the overall quality of audio samples. Likewise, we assess the ability of audio scriptwriting, rated from one to fice, from five different aspects: audio-text relevance, audio coherence, naturalness, engagement, and creativity. To extend AUDIT to the audio scriptwriting task, we use chatgpt to convert the instruction to several basic tasks and call AUDIT to execute tasks recursively. We refer to the extended AUDIT AUDIT$+$ in this paper.